This package uses the natbib.sty package to generate author-year citations and bibliographies. There is much more you can do than what’s on this page. Search the web and your TeX system for “natbib.dvi” and “natbib.sty” for clues.

To get a citation in the main text, read Ahrens et al. (1994a). You can get several letters on a year, or several years, if you follow Ahrens et al. (1994a,b). Watch out, however, and avoid messes like Ahrens et al. (1994a), Hammel et al. (1995), since you would want an an “and” before the final citation. Just list such citations separately like Ahrens et al. (1994a) and Hammel et al. (1995).

Parenthetical citations also work (McGregor et al. 1996), and you can string as many as you like together (Hammel et al. 1995, McGregor et al. 1996, Noll et al. 1996) was made.

There are many other things you can do (Hammel et al. 1995, hereafter HAM95; Harrington et al. 1994; McGregor et al. 1996; Noll et al. 1996). You can even put notes on a parenthetical citation (the book by Noll et al. 1996 is excellent). If you want just the author’s name, follow Zahnle and Mac Low.

Here’s how to get a references-style entry anywhere you like, say for a CV (remember that you have to add the final period back in): Ahrens, T. J., T. Takata, J. D. O’Keefe, and G. S. Orton 1994a. Impact of

comet Shoemaker-Levy 9 on Jupiter. Geophys. Res. Lett. 21, 1087–1090.

References

Ahrens, T. J., T. Takata, J. D. O’Keefe, and G. S. Orton 1994a. Impact of comet Shoemaker-Levy 9 on Jupiter. Geophys. Res. Lett. 21, 1087–1090.

Ahrens, T. J., T. Takata, J. D. O’Keefe, and G. S. Orton 1994b. Radiative signatures from impact of comet shoemaker-levy-9 on jupiter. Geophys. Res. Lett. 21, 1551–1553.

Hammel, H. B., R. F. Beebe, A. P. Ingersoll, G. S. Orton, J. R. Mills, A. A. Simon, P. Chodas, J. T. Clarke, E. de Jong, T. E. Dowling, J. Harrington, L. F. Huber, E. Karkoschka, C. M. Santori, A. Tiogo, D. Yeomans, and R. A. West 1995. Hst imaging of atmospheric phenomena created by the impact of comet shoemaker-levy 9. Science 267, 1288–1296.

Harrington, J., R. P. Le Beau, K. A. Backes, and T. E. Dowling 1994. Dynamic response of jupiter’s atmosphere to the impact of comet shoemaker-levy-9. Nature 368, 525–527.

McGregor, P. J., P. D. Nicholson, and M. G. Allen 1996. Caspir observations of the collision of comet shoemaker-levy 9 with jupiter. Icarus 121, 361–388.

Noll, K. S., H. A. Weaver, and P. D. Feldman (Eds.) 1996. The Collision of Comet Shoemaker-Levy 9 and Jupiter, IAU Colloquium 156, Cambridge, UK. Cambridge University Press.

Zahnle, K., and M.-M. Mac Low 1995. A simple model for the light curve generated by a shoemaker-levy 9 impact. J. Geophys. Res. 100, 16885–16894.